OpenStudy (anonymous):

Simplify (x-4) (x^2+3x-1) please I need help

OpenStudy (anonymous):

what i got was x^3-x^2-13x+4

OpenStudy (anonymous):

i did that by multiplying everything in (x^2 + 3x -1) by x and then by 4 and combined like terms

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok ARoC (average rate of change) = change in y over change in x. so from -3 to 5, you y changed by 8 and from -3 to 1, your x changed by 4 so 8/4 = 2. your ARoC is 2

OpenStudy (anonymous):

take (3xy-18y) + (2x-12) take out 3y in the first part so you have 3y(x-6) and then take out 2 in the other part so you have 2(x-6). so. 3y(x-6)+2(x-6)

OpenStudy (anonymous):

so then in the end you should get (x-6)(3y+2)

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

we get the same thing we just got it different ways

OpenStudy (anonymous):

course :)

OpenStudy (anonymous):

yes, it is

OpenStudy (anonymous):

ill check it, gimme a sec.

OpenStudy (anonymous):

Yes, 43 is correct

OpenStudy (anonymous):

yes. so first i found t(4) = 4(4)+3 so t(4)= 19. next i found c(19) = 2(19) + 5 = 43

OpenStudy (anonymous):

( thank you for your testimonial :) im glad to know im helping)

OpenStudy (anonymous):

im happy to help whenever :)

OpenStudy (anonymous):

im figuring out the how but your answer should be\[4\sqrt[4]5{}\]

OpenStudy (anonymous):

https://www.desmos.com/calculator/2npgb907ik

OpenStudy (anonymous):

C is your answer

OpenStudy (anonymous):

yes, youre right

OpenStudy (anonymous):

if you change c into y=mx+b form, you get y-3=2x-4. add the 3, y=2x-1 which is the same as the line i first gave you :)

OpenStudy (anonymous):

ok so before you even plug in 3, find the inverse of your original equation. do you know how to do that?

OpenStudy (anonymous):

im not sure if you responded...my computer hates me XD

OpenStudy (anonymous):

if you didnt, then the way you find it is you basically realize that f(x) is a fancy way of saying y, so you switch the x and the y in your problem and resolve again for y. so x = (y-5)/3

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

your inverse function is 3x+5. now substitute 3 in for x and you get 9+5 which is 14 so yes you got it correct, good job

OpenStudy (anonymous):

sorry, had to go take food out of the oven before my house burns to the ground XD

OpenStudy (anonymous):

worthy cause XD

OpenStudy (anonymous):

Yes, your answer is B

OpenStudy (anonymous):

ok. the conjugate of a complex number is a long and complicated way of saying leave the real number alone, change the sign of the imaginary number

OpenStudy (anonymous):

so since 2 is your real and -8i is your imaginary, change the - to a +8i and you get 2+8i

OpenStudy (anonymous):

here, this explains it http://www.mathwords.com/c/complex_conjugate.htm

OpenStudy (anonymous):

awesome :)

OpenStudy (anonymous):

anything else you need help with?

OpenStudy (anonymous):

i got all the time you need :)

OpenStudy (anonymous):

sure thing. :) youre talking to an insomniac, i got a lot of time XD

OpenStudy (anonymous):

do you know the x factor? it makes things pretty simple to at least see

OpenStudy (anonymous):



OpenStudy (anonymous):

in an equation where there is a number before your x^2, multiply that number by your last number

OpenStudy (anonymous):

so you have \ 68 / \ / \ / \/ / \ / \ / \ / 16 \ just pretend that looks decent for me XD

OpenStudy (anonymous):

now you find what number multiplies to make 68 and adds to make 16

OpenStudy (anonymous):

D is your answer

OpenStudy (anonymous):

im just confusing you XD sorry.

OpenStudy (anonymous):

haha thank you

OpenStudy (anonymous):

A. 2. im figuring out how to explain that though....

OpenStudy (anonymous):

ill check right now

OpenStudy (anonymous):

here, @awkwardpanda will also help too :)

OpenStudy (anonymous):

ya, but im gonna be eating at the same time and i dont want my answers to get too delayed XD

OpenStudy (anonymous):

there they are :)

OpenStudy (anonymous):

ok for that one, you could do solving using the quadratic formula. but for me its easier to subtract 20 and have 2x^2+3x -20 = 0

OpenStudy (anonymous):

so what multiplies to -40 and adds to 3?

OpenStudy (anonymous):

yes good job :)

OpenStudy (anonymous):

im sorry, i was eating

OpenStudy (anonymous):

ok this one might take me a little so hold on a sec

OpenStudy (anonymous):

im lazy so im just gonna plug the #s in XD

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

anything else ya need?

OpenStudy (anonymous):

sure thing. dont really got anything better to do XD

OpenStudy (anonymous):

sorry about this but your messages (if there are any) arent coming in all that well on my comp. so i cant tell if youre saying anything or not...

OpenStudy (anonymous):

ok now this one goes y=50. y=10x and y=5x+20

OpenStudy (anonymous):

do you see why?

OpenStudy (anonymous):

anything else?

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

yes, they both started with the same expression. figuring out how they got their answers

OpenStudy (anonymous):

they both started with x^(1/3)

OpenStudy (anonymous):

im not too sure about this one, sorry.

OpenStudy (anonymous):

yeah, im not the best at word problems XD

OpenStudy (anonymous):

any parabola such as y=x^2 would work, the focus of y=x^2 is at (0,1/4) and the directrix y=1/4

OpenStudy (anonymous):

cool :)